OpenStudy (anonymous):

ABSTRACT ALGEBRA: pROVE THAT THE SET OF ALL 2X2 MATRICES WITH ENTRIES FROM r AND DETERMINANT +1 IS A GROUP UNDER MATRIX MULTIPLICATION. @ash2326 @robtobey @whpalmer4 @wio @Compassionate @shamil98 @beccaboo333 @nincompoop @wolfe8 @Zale101 @Ashleyisakitty @e.cociuba @Isaiah.Feynman @kewlgeek555 @tester97 @Gabylovesyou

OpenStudy (beccaboo333):

Sorry I'm not good with math.

OpenStudy (tester97):

Sorry im barely passing alg 2 i doubt i could help you with that.

OpenStudy (science0229):

First, let's let G be all of the 2x2 matrices

OpenStudy (science0229):

Now, do you know the definition of a group under matrix multiplication?

OpenStudy (anonymous):

Yep! It has to have closure, associativity, an identity element, and an inverse! (Took linear algebra last semester so I know all these cool terms lol)

OpenStudy (science0229):

So our goal is then to prove all of them.

OpenStudy (science0229):

Let's start with the easy one.

OpenStudy (anonymous):

Can you please tell me what the notation GL(2,R) is in abstract algebra?

OpenStudy (science0229):

That means the set of all invertible matrices in Real number

OpenStudy (science0229):

Correction: the set of 2x2 invertible matrices over real number

OpenStudy (science0229):

Now, one of the definitions was that it has to include the identity matrix, right?

OpenStudy (anonymous):

@science0229 Now what does SL(2,R) mean then?

OpenStudy (science0229):

SL(2,R) is the set of 2x2 matrices with determinant of 1 over real number

OpenStudy (anonymous):

How do you know that? Haha

OpenStudy (science0229):

I know many random things. I was studying matrices. I got curious. I went in a little bit deep, and that came up...

OpenStudy (anonymous):

@science0229 How is {1,2,3,4} under multiplication modulo 5 a group?

OpenStudy (science0229):

Hey. I don't know that deep! I'm just an 8th grader with curiosity.

OpenStudy (anonymous):

Here how you do it 1) closure det (A B)= det(A) det(B) =1x1 =1 2) If A is in G, then the inverse of A is \[ det(A^{-1}) = \frac 1 {det(A)}= \frac 1 1=1 \]

OpenStudy (anonymous):

The identity matrix I belong to G since det(I) =1 That is all what you need to show. Since Multiplication of matrices is associative